Find four consecutive positive integers such that the product of the first and the fourth is four less than twice the first multiplied by the fourth
Answer by mananth(16946)

let the integers be x, x+1, x+2, x+3
x(x+3)=2x(x+3)-4
x^2+3x=2x^2+6x-4
x^2+3x-4=0
x^2+4x-x-4=0
x(x+4)-1(x+4)=0
(x+4)(x-1)=0
x=1 which is positive
1,2,3,4